An airplane trip involves three legs, with two stopovers. The first leg is due East for 620 km the second leg is Southeast (45
1
Expert's answer
2021-10-01T12:22:41-0400

Although the condition is incomplete, assume the second leg is 45° SE and it is 500 km long. The displacement between the airport of departure and the first stopover is


"D_E=620+500\cos45°=974\text{ km}.\\
D_S=0+500\sin45°=354\text{ km},\\\space\\
D=\sqrt{D_E^2+D_S^2}=1036\text{ km}."